Calculate the formula mass of(a.m.u) of the AL2o3



1
Expert's answer
2021-10-31T12:40:56-0400

"= 26.9815 \times 2 + 15.9994 \times 3 \\

= 53.963 + 47.9982 \\

= 101.9612 \; a.m.u."